OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (4, -2) and perpendicular to y = x - 7.

OpenStudy (ghazi):

line perpendicular to the given line will have slope = negative of reverse of the slope of given line= -1 now the required equation of line will be y=-x+d ...now to find d substitute the points (4,-2) in the line y=-x+d at the place of x and y ....you'll get d= y+x=-2+4=2 therefore your final equation of line will be y=-x+2

OpenStudy (anonymous):

thanks :)! can you help me with more @ghazi

OpenStudy (ghazi):

sure !!

OpenStudy (anonymous):

4_01]Which of the following is the slope between the points (-6, 0) and (2, -9)? - -

OpenStudy (anonymous):

8/9 -8/9 9/8 -9/8

OpenStudy (ghazi):

use (y2-y1)/(x2-x1).....(-9-0)/(2-(-6))

OpenStudy (anonymous):

those arethe choices and ok hmm

OpenStudy (ghazi):

what you got?

OpenStudy (anonymous):

c you?

OpenStudy (ghazi):

it's d....

OpenStudy (anonymous):

i mustve done something wrong

OpenStudy (anonymous):

4_07] (-1, 3) and perpendicular to y = -1/3x + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

i got b

OpenStudy (ghazi):

y=3x+d...d= 6....therefore it's b

OpenStudy (anonymous):

yesa!!!

OpenStudy (anonymous):

i got it !

OpenStudy (anonymous):

4_05]Choose the equation of the horizontal line that passes through the point (2, 3). x = 2 x = 3 y = 2 y = 3

OpenStudy (anonymous):

a

OpenStudy (ghazi):

it's y=3

OpenStudy (anonymous):

can you explain that one to me if you can

OpenStudy (ghazi):

just y=3 is a horizontal line that is parallel to x axis and passes through 3

OpenStudy (anonymous):

ohhhh

OpenStudy (ghazi):

 at every value of x y=3 so at x=2 again y=3

OpenStudy (anonymous):

Identify the type of correlation in the graph below. Negative Correlation Positive Correlation No Correlation Not Enough Information

OpenStudy (anonymous):

ohhi see! thanks :) can you help me with ^

OpenStudy (ghazi):

negative

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (ghazi):

d

OpenStudy (anonymous):

Which of the following is an example of positive correlation? House address vs math test score Age vs commute time to school Time worked vs money earned Time eating vs food left

OpenStudy (anonymous):

this one is really weird

OpenStudy (ghazi):

c..the more you work the more you'll be paid

OpenStudy (ghazi):

hope you're done now :))

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1) aha:) i have sme more :)

OpenStudy (ghazi):

b

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = x + 2. y = 4x - 1 y = 4x - 3 y = x 1/4- 3 y = x1/4 - 1

OpenStudy (ghazi):

none of the options are correct

OpenStudy (anonymous):

theres no way oh sorry i forog t

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/3x + 2. y = 4x - 1 y = 4x - 3 y = x 1/4- 3 y = x1/4 - 1

OpenStudy (ghazi):

he he...it's c..but still i have doubt

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = 1/3x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (ghazi):

c

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.